Assume you are making business decision on alternative two businesses under determined
risks. Given the following probability of the risks and estimated profit, (5 marks)
a. Calculate the expected profit you could get under each risk probabilities for the
business alternatives.
b. Which business do you select based on the expected profit?
c. By constructing a discrete probability distribution, which business do you select?
d. By calculating the expected mean values, variance, standard deviation and
coefficient of variation, which business do you select?
Business Economic Condition Probability Outcome/
Profit
Expected
Profit
Y
Good economic condition/boom 0.2 900
Normal economic condition 0.5 800
Bad economic condition/recession 0.3 700
Expected Profit
Z
Good economic condition/boom 0.2 1000
Normal economic condition 0.5 800
Bad economic condition/recession 0.3 600
Expected Profit�
1
Expert's answer
2021-08-23T13:11:43-0400

a)expected profit=good condition-normal condition

"y=0.58-0.29=0.29"

"z=0.58-0.21=0.37"

b)z because it has high expected profit

c)

y

Good economic condition/boom 0.2 900

Normal economic condition 0.5 800

Bad economic condition/recession 0.3 700

expected profit 0.29

Z

Good economic condition/boom 0.2 1000

Normal economic condition 0.5 800

Bad economic condition/recession 0.3 600

expected profit 0.37

z because it has high profit and the probability of recession is also low.

d)

"y=\frac{0.29+0.58+0.37}{3}=0.99"

"z=\frac{0.21+0.58+0.36}{3}=0.38"

y it has a high probability of happening